OpenStudy (anonymous):

( 9 / (x+1) ) - (1/2) = (2/ (3x+3) )

OpenStudy (anonymous):

x=10/3

OpenStudy (anonymous):

incorrect?

OpenStudy (anonymous):

I attached the problem.

OpenStudy (anonymous):

x = 47/3 or 15.6666... Multiply all terms by (x+1) 9 - 2/3 = 1/2 * (x+1) 18-4/3 = x+1 x = 16 -1/3